OpenStudy (anonymous):

Need some help plez two planes leave o'hare airport in chicago one flies directly east at 540 miles and leaves at 1pm the other leaves directly west at 720 miles per hour at what time will they be 5130 miles apart

OpenStudy (anonymous):

Do both planes leave at 1pm ?

OpenStudy (anonymous):

the 1st leaves at 1pm the leaves at noon

OpenStudy (anonymous):

r t = d, t = d/r\[\frac{5130-720}{540+720}=\frac{7}{2} \]1 PM + 3.5 hours = 4:30 PM

OpenStudy (anonymous):

thank youu soo much!!!!!

OpenStudy (anonymous):

Your welcome.